OpenStudy (fran345):

I need help. I don't know what I'm doing. 1.) 2x + y = 3 2.) x - 2y = -1 If equation 1 is multiplied by 2 and then the equations are added, the result is A.3x=2 B.3x=5 C.5x=5

OpenStudy (poison_princess1):

Here download this app it very helpful for problems like this: Photomath

OpenStudy (poison_princess1):

https://photomath.net/en/

OpenStudy (3mar):

Well, I am here. Could I help you?

OpenStudy (poison_princess1):

Hi

OpenStudy (fran345):

yes

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Where are you stuck?

OpenStudy (fran345):

from the start

OpenStudy (poison_princess1):

Here the steps for the first problem https://photomath.net/s/M08Vj Second part https://photomath.net/s/B7RV0

OpenStudy (poison_princess1):

The do the rest on ur own Hope it helps

OpenStudy (3mar):

\[\huge 2x + y = 3\] ...... 1 \[\huge x - 2y = -1\] ...... 2 multiply 1 by 2 and then add to 2, you get \[\huge 4x + 2y = 6\] \[\huge x - 2y = -1\] ---------------------- \[\huge5x=5\] Got it?

OpenStudy (fran345):

yes, can you help with another one?

OpenStudy (3mar):

Of course.